OpenStudy (zzr0ck3r):

equilibrium solutions to system of equations dr/dt=2*R*(1- R/2) - 1.2*R*F dF/dt=-F + 0.9*R*F

OpenStudy (experimentx):

??? the system is in equilibrium when, dF/dr = 0

OpenStudy (zzr0ck3r):

yes it is

OpenStudy (experimentx):

dF/dr = dF/dt * dt/dr = (-F + 0.9*R*F)/(2*R*(1- R/2) - 1.2*R*F) = 0

OpenStudy (zzr0ck3r):

this is where I'm not sure how to do this. I have three points but I dont know if that is all

OpenStudy (experimentx):

(-F + 0.9*R*F) = 0 R = 1/0.9

OpenStudy (zzr0ck3r):

yes I have (0,0), (2,0), and (10/9,20/27)

OpenStudy (zzr0ck3r):

I am just doing this by intuition, is there some more systematic way of doing it?

OpenStudy (experimentx):

dF/dR = (-F + 0.9*R*F)/(2*R*(1- R/2) - 1.2*R*F) equate it to zero , (-F + 0.9*R*F)/(2*R*(1- R/2) - 1.2*R*F) = 0 and solve for R, answer is, R = 10/9 To find F, integrate or solve this differential equation dF/dr = (-F + 0.9*R*F)/(2*R*(1- R/2) - 1.2*R*F) F = F(R), put the value of R=10/9 you should get 20/27

OpenStudy (zzr0ck3r):

so is there only 3 equilibrium points?

OpenStudy (zzr0ck3r):

because (0,0) and (2,0) are solutions alos

OpenStudy (experimentx):

not sure ... but this must be sure for one, r = 10/9

OpenStudy (zzr0ck3r):

hmm yeah I need all, and many times there are more than one. (0,0) and (2,0) you can see from looking at it but some are more dificult to get to.

OpenStudy (zzr0ck3r):

yeah haha tnx m8